# DeepQuestion: Systematic Generation of Real-World Challenges for Evaluating LLMs Performance

### v1

@@ Appendix B Prompts
Deep-Question Prompt in Q2I for Physics Questions My goal is to design questions based on Bloom’s *Assessment* level. I will provide you with a question. Your task is to write a simple, non-detailed instruction that students will use to design a new question**—one that leads to **exactly the same final answer as the original question. The aim is to assess whether students have a deep enough understanding of the concept to create a question that results in a specific outcome. ### The following conditions must be met: * The instruction should not explicitly suggest the use of any formula. * The instruction should not include specific details about question design (e.g., what parameters or values are needed). * The final output (i.e., the answer) must be clearly stated in the instruction and must match the original question’s final answer. * Since the instruction is given independently of the original question, it must include all the context/data required. * The instruction should not provide direct guidance and must resemble an exam-style prompt. * The designed question must not be multi-part; it should have one clear requirement that is also explicitly stated. * All formulas and numbers must be presented in LaTeX format. * The instruction must be written in Persian. ### Example: Design a question that, using the concepts of motion, results in a space-time equation with the final output $x=3t-4$
Prompt for Evaluation of Answers You are a helpful assistant. You will receive a physics question. Your task is to answer it in Persian using English numerals and LaTeX format for all formulas and numerical expressions. Solve the problem step by step (chain-of-thought reasoning) before providing the final answer. Ensure clarity and logical reasoning throughout your explanation. If the final answer is a number or formula, repeat it exactly at the end, after the following line: ### answer: Only the final answer should appear after the "### answer:" line, with no extra text, punctuation, or explanation. Do not include anything after the "### answer:" section.
System Prompt Question Generator in Q2I "You are a helpful assistant. Design a physics question based on the input instruction and consider the following: * All formulas and numbers should be in English and LaTeX format. * All your answers must be in Persian. * Just design the question; do not answer it. * Do not include the final answer in your question. * The designed question should not be multi-part; it should ultimately have one requirement, and that requirement must also be specified in the instruction given to you. * Remember you are NOT ALLOWED to give the final answer explicitly in the question. * You must think step by step before generating the question. Put your reasoning process between <think> and </think> tags. * The final designed question in Persian must be written immediately after the </think> tag."
System Prompt for Prompt Generator in Physics Questions You are a prompt engineer tasked with writing a Scenario Generation Instruction (SGI). This SGI will be used to guide a language model in generating original, scenario-based physics questions. Your SGI must instruct the model to: Create questions grounded in realistic or imaginative physical scenarios. Ensure each question tests the same underlying physics principles as a provided set of example questions. Avoid copying or making superficial edits to the example questions. Your final output should be a well-structured, standalone prompt that clearly communicates these instructions to the model. Output only the SGI.
System Prompt for Prompt Evaluator in Physics Questions You are a helpful assistant assigned to evaluate a Scenario Generation Instruction (SGI), which is designed to guide another model in generating high-quality, scenario-based physics questions based on a sample set of physics Q&A. Your task is to: Evaluate the SGI according to these criteria: Clarity: Are the instructions understandable and unambiguous? Relevance: Does it stay focused on the concepts represented in the example questions? Instructional Effectiveness: Will it likely result in creative and accurate scenario-based questions? Creative Direction: Does it provide sufficient guidance to inspire realistic or imaginative scenarios rooted in physics? Provide Feedback: Suggest specific improvements to strengthen the SGI. Score the SGI: Assign a numerical score between 00 (very poor) and $10$ (excellent) based on overall quality. Your output must include: A brief evaluation summary (strengths and weaknesses). Clear, actionable suggestions for improvement. A final numeric score between 00 and $10$.
Appendix C Q2S Results
These charts show each model’s accuracy on the original benchmark questions compared to their scenario-based (Q2S) versions, illustrating the performance drop when questions are embedded in realistic narratives.
Figure 4: Result of Original vs.Q2S in Physics
Figure 5: Result of Original vs. Q2S in GSM8K
Appendix D Quality of Generated Question
These plots display, for each of five quality criteria, the percentage of cases where original questions “win” over generated ones, highlighting which aspects (e.g., reasoning demand, clarity) most favor the originals.
Figure 6: Win rate original against Q2I for O4-mini in GSM8K questions
Figure 7: Win rate original against Q2I for Deepseek-R1 in physics questions
Figure 8: Win rate original against Q2I for Deepseek-R1 in GSM8K questions
Figure 9: Win rate original against Q2I for Qwen3-32b in physics questions
Figure 10: Win rate original against Q2I for Qwen3-32b in GSM8K questions

### v2

@@ 9 $(score,feedback)\leftarrow\texttt{LLM}(evaluation\_criteria,prompt)$;
(a) Original vs. Q2S in GSM8K
(b) Original vs. Q2S in Physics
Figure 4: Evaluation of different LLMs in original and scenario-based questions
(a) Original vs. Q2I in GSM8K
(b) Original vs. Q2I in Physics
Figure 5: Evaluation of different LLMs in original and instruction-based questions
4.  Experiment Setup
(a) GSM8K
(b) Physics
Figure 6: Win rate of original against Q2I questions for O4-mini
(a) GSM8K
(b) Physics
Figure 7: Win rate of original against Q2I questions for Deepseek-R1
(a) GSM8K
(b) Physics
Figure 8: Win rate of original against Q2I questions for Qwen3-32b
In the Q2I setting, the model generates questions from instructions. Evaluation consisted of two steps: (1) Answerability Check: verifying whether a strong model (O4-mini) and the model itself can correctly solve the generated questions, indicating questions validity. (2) LLM-as-Judge: since answerability alone does not capture question quality, a powerful language model (O4-mini) was employed for direct qualitative evaluation based on expert-defined criteria: Reasoning Demand — how much genuine thinking and decision-making the question requires; Numerical Quality — whether the numerical values are realistic and meaningful; Physical Realism — whether the scenario is plausible and internally consistent; Clarity and Brevity — how concise and understandable the question is; and Solution Spoiling — whether the problem avoids revealing its own answer or solution steps.
In future work, for open-ended questions that do not have a single correct answer, the same LLM-as-Judge approach could be extended. By defining domain-specific evaluation criteria, LLMs can provide consistent, qualitative assessments that capture nuanced aspects of question quality beyond simple answerability.
Table 1: Model performance across translation tasks.
(a) Physics – Translate to English
Models Llama3-3-70b Llama3-3-70b-Trans GPT-4.1 GPT-4.1-Trans O4-mini O4-mini-Trans
Original Questions 48.33% 61.67% 76.67% 78.33% 83.33% 78.33%
Q2S 45.00% 50.00% 71.67% 75.00% 76.67% 80.00%
Q2I 18.33% 23.33% 26.67% 33.33% 80.00% 73.33%
(b) GSM8K – Translate to Persian
Models Llama3-3-70b Llama3-3-70b-Trans GPT-4.1 GPT-4.1-Trans O4-mini O4-mini-Trans
Original Questions 96.67% 88.33% 98.33% 96.67% 100.00% 100.00%
Q2S 96.67% 75.00% 100.00% 95.00% 100.00% 98.33%
Q2I 25.00% 11.67% 36.67% 31.67% 96.67% 95.00%
5.  Results
Based on the distinctions between general-purpose models and reasoning-focused models, we evaluated LLMs from both categories and report the results. Our evaluation methodology is similar to that of Illusion of Thinking, which also compared these two families of models.
5.1.  General-Purpose LLMs
First, we evaluated general-purpose LLMs on question generation using Q2S. As shown in Figure˜4, general-purpose models exhibit a moderate decrease in performance on Q2S questions. Some models, such as Gemini-2-Flash and LLaMA-3.1-8B, demonstrate a noticeable reduction in accuracy, whereas others, including Gemma-3-27B and DeepSeek-v3, achieve comparable or even improved performance on Q2S questions. These findings indicate that general-purpose LLMs can solve scenario-based questions with only a slight decrease in accuracy, suggesting their capability to apply knowledge at the applying level of Bloom’s taxonomy.
To further assess LLMs’ abilities at higher levels, such as the creation level, we were motivated to design the Q2I experiment. In this setting, models exhibited a substantial drop in accuracy when tackling tasks that require deeper understanding and more complex reasoning. No model surpassed 38% accuracy on instruction-based question generation for GSM8K or physics (see Figure˜5), despite over 95% on original questions. These results support the use of Bloom’s taxonomy as an evaluation framework and reveal the limitations of general-purpose LLMs on higher-order cognitive tasks.
5.2.  Reasoning LLMs
Based on the observed results from general-purpose models and their performance decline on Q2I questions, we evaluated reasoning models on both Q2I and Q2S questions. Reasoning models exhibited a performance trend similar to that of general-purpose models on Q2S questions. However, for Q2I questions, the performance reduction in reasoning models was smaller. Despite this, accuracy in the Q2I setting still decreased by 8–30% (Figure˜5), highlighting limitations of these models at higher cognitive levels. For example, Deepseek-R1’s performance declined by 30% on GSM8K and 10% on physics datasets, whereas Qwen-3-32B experienced an 11% drop on both datasets. Although reasoning models generally outperform general-purpose models, significant gaps remain in achieving deep and comprehensive understanding. These findings indicate that neither general-purpose nor reasoning models achieve strong performance at the creation level of Bloom’s taxonomy, suggesting that current LLMs are still distant from high-level cognitive reasoning.
5.3.  Question Quality Analysis
While reasoning models can generate questions that follow instructions and maintain topical consistency, it is crucial to assess their quality beyond surface-level compliance. To this end, we use a preference-based evaluation with O4-mini as a judge, directly comparing generated and original questions across five criteria: reasoning demand, numerical quality, physical realism, clarity and brevity, and solution spoiling. This approach addresses concerns about the model’s ability to assess quality by focusing on direct comparison.
As illustrated in Figures˜6, 7 and 8, the physics questions generated by the models continue to underperform compared to the original questions. Specifically, in the dimensions of Reasoning Demand, Creativity, and Brevity, the original questions outperform the generated ones in more than 50% of cases across all three models. Furthermore, in terms of Solution Spoiling, Q2I exhibits negligible improvement, achieving wins in only approximately 3% of cases. These findings underscore the persistent challenges faced by reasoning models in producing high-quality questions. In contrast, generated GSM8K questions are more comparable to the originals, likely reflecting the relative simplicity of this dataset. Overall, the results demonstrate the effectiveness of our framework in enhancing question depth without introducing artificial difficulty. Nevertheless, a substantial quality gap remains, emphasizing the ongoing difficulty of generating high-quality questions even with advanced reasoning models.
5.4.  Disentangling Linguistic Effects from Cognitive Complexity
To rule out linguistic artifacts, we conducted a cross-lingual experiment by translating the Persian Physics dataset into English (see Table˜1(a) and the English GSM8K dataset into Persian (see Table˜1(b)). We evaluated three representative models, including Llama3-70B, GPT-4.1, and O4-mini, to test whether performance declines in DeepQuestion stem from increased cognitive demand or from language-related confounds.
Across both datasets and translation directions, the performance hierarchy (Original > Q2S > Q2I) remained consistent. For instance, in the translated Physics dataset, O4-mini scored 83.33% (Original), 80.00% (Q2S), and 73.33% (Q2I); GPT-4.1 followed the same pattern (78.33% → 75.00% → 33.33%). Similarly, in Persian-translated GSM8K, Llama3-70B dropped from 88.33% to 75.00% and then to 11.67%. This stability confirms that our transformations genuinely elevate cognitive complexity independent of language.
These results confirm that DeepQuestion isolates cognitive complexity as defined by Bloom’s Taxonomy. Observed declines arise from reasoning demands rather than language artifacts, validating our framework’s robustness across languages and model families.
6.  Conclusion
This work presented DeepQuestion, a systematic and scalable framework for generating cognitively diverse benchmarks that extend existing datasets through the lens of Bloom’s taxonomy. By transforming conventional question–answer pairs into scenario-based and instruction-driven tasks, DeepQuestion provides a structured means to probe LLMs across multiple levels of cognition, from application to creation.
Our findings across state-of-the-art LLMs reveal a persistent decline in performance as task complexity and cognitive depth increase—reaching up to 70% accuracy loss in higher-order reasoning tasks. These results highlight that current LLMs, despite their success on standardized benchmarks, still exhibit shallow generalization and limited conceptual transfer when confronted with real-world or creative problem-solving scenarios.
Beyond performance assessment, DeepQuestion offers a replicable pathway for future benchmark development. The framework can be readily applied to other domains such as law, medicine, or engineering to examine domain-specific reasoning. Moreover, it opens new directions for exploring automated benchmark construction, self-improving evaluation pipelines, and curriculum-style training data aligned with cognitive learning theories.
In summary, DeepQuestion bridges educational psychology and AI evaluation, revealing fundamental gaps in LLM reasoning and offering tools to measure and close them. As language models continue to evolve, cognitively grounded and contextually rich evaluation frameworks like DeepQuestion will be essential for steering their progress toward genuine understanding and human-aligned intelligence.
7.  Bibliographical References
[…]
  • Abdin et al. (2024) Marah Abdin, Jyoti Aneja, Harkirat Behl, Sébastien Bubeck, Ronen Eldan, Suriya Gunasekar, Michael Harrison, Russell J Hewett, Mojan Javaheripi, Piero Kauffmann, et al. 2024. Phi-4 technical report. arXiv preprint arXiv:2412.08905.
  • Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al. 2023. GPT-4 technical report. arXiv preprint arXiv:2303.08774.
  • Alaa et al. (2025) Ahmed Alaa, Thomas Hartvigsen, Niloufar Golchini, Shiladitya Dutta, Frances Dean, Inioluwa Deborah Raji, and Travis Zack. 2025. Medical large language model benchmarks should prioritize construct validity. arXiv preprint arXiv:2503.10694.
  • Anderson and Krathwohl (2001) Lorin W Anderson and David R Krathwohl. 2001. A taxonomy for learning, teaching, and assessing: A revision of Bloom’s taxonomy of educational objectives: complete edition. Addison Wesley Longman, Inc.
  • Arias-Duart et al. (2025) Anna Arias-Duart, Pablo Agustin Martin-Torres, Daniel Hinjos, Pablo Bernabeu-Perez, Lucia Urcelay Ganzabal, Marta Gonzalez Mallo, Ashwin Kumar Gururajan, Enrique Lopez-Cuena, Sergio Alvarez-Napagao, and Dario Garcia-Gasulla. 2025. Automatic evaluation of healthcare llms beyond question-answering. arXiv preprint arXiv:2502.06666.
  • Biggs and Collis (2014) John B Biggs and Kevin F Collis. 2014. Evaluating the quality of learning: The SOLO taxonomy (Structure of the Observed Learning Outcome). Academic press.
  • Bloom (1971) Benjamin S Bloom. 1971. Taxonomy of educational objectives: The classification of educational goals: By a committee of college and university examiners. David McKay.
  • Chauhan et al. (2025) Archana Chauhan, Farah Khaliq, and Kirtana Raghurama Nayak. 2025. Assessing quality of scenario-based multiple-choice questions in physiology: Faculty-generated vs. chatgpt-generated questions among phase i medical students. International Journal of Artificial Intelligence in Education.
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training verifiers to solve math word problems. arXiv preprint arXiv:2110.14168.
  • Forehand (2010) Mary Forehand. 2010. Bloom’s taxonomy. Emerging perspectives on learning, teaching, and technology, 41(4).
  • Grattafiori et al. (2024) Aaron Grattafiori, Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Alex Vaughan, et al. 2024. The llama 3 herd of models. arXiv preprint arXiv:2407.21783.
  • Guo et al. (2025) Daya Guo, Dejian Yang, Haowei Zhang, Junxiao Song, Ruoyu Zhang, Runxin Xu, Qihao Zhu, Shirong Ma, Peiyi Wang, Xiao Bi, et al. 2025. Deepseek-r1: Incentivizing reasoning capability in llms via reinforcement learning. arXiv preprint arXiv:2501.12948.
  • Hendrycks et al. (2021a) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2021a. Measuring massive multitask language understanding. Proceedings of the International Conference on Learning Representations.
  • Hendrycks et al. (2021b) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt. 2021b. Measuring mathematical problem solving with the math dataset. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2).
  • İlhan et al. (2017) Mustafa İlhan, Melehat Gezer, et al. 2017. A comparison of the reliability of the solo-and revised bloom’s taxonomy-based classifications in the analysis of the cognitive levels of assessment questions. Pegem Egitim ve Ogretim Dergisi= Pegem Journal of Education and Instruction, 7(4):637.
  • Leung (2000) CF Leung. 2000. Assessment for learning: Using solo taxonomy to measure design performance of design & technology students. International Journal of Technology and Design Education, 10(2).
  • Liu et al. (2024) Aixin Liu, Bei Feng, Bing Xue, Bingxuan Wang, Bochao Wu, Chengda Lu, Chenggang Zhao, Chengqi Deng, Chenyu Zhang, Chong Ruan, et al. 2024. Deepseek-v3 technical report. arXiv preprint arXiv:2412.19437.
  • Mirzadeh et al. (2025) Seyed Iman Mirzadeh, Keivan Alizadeh, Hooman Shahrokhi, Oncel Tuzel, Samy Bengio, and Mehrdad Farajtabar. 2025. Gsm-symbolic: Understanding the limitations of mathematical reasoning in large language models. In The Thirteenth International Conference on Learning Representations.
  • Myrzakhan et al. (2024) Aidar Myrzakhan, Sondos Mahmoud Bsharat, and Zhiqiang Shen. 2024. Open-llm-leaderboard: From multi-choice to open-style questions for llms evaluation, benchmark, and arena. arXiv preprint arXiv:2406.07545.
  • Petrov et al. (2025) Ivo Petrov, Jasper Dekoninck, Lyuben Baltadzhiev, Maria Drencheva, Kristian Minchev, Mislav Balunović, Nikola Jovanović, and Martin Vechev. 2025. Proof or bluff? evaluating llms on 2025 USA math olympiad. arXiv preprint arXiv:2503.21934.
  • Rein et al. (2024) David Rein, Betty Li Hou, Asa Cooper Stickland, Jackson Petty, Richard Yuanzhe Pang, Julien Dirani, Julian Michael, and Samuel R Bowman. 2024. GPQA : A graduate-level google-proof q&a benchmark. In First Conference on Language Modeling.
  • Shojaee et al. (2025) Parshin Shojaee, Iman Mirzadeh, Keivan Alizadeh, Maxwell Horton, Samy Bengio, and Mehrdad Farajtabar. 2025. The illusion of thinking: Understanding the strengths and limitations of reasoning models via the lens of problem complexity. arXiv preprint arXiv:2506.06941.
  • Team et al. (2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, Katie Millican, et al. 2023. Gemini: a family of highly capable multimodal models. arXiv preprint arXiv:2312.11805.
  • Team et al. (2025) Gemma Team, Aishwarya Kamath, Johan Ferret, Shreya Pathak, Nino Vieillard, Ramona Merhej, Sarah Perrin, Tatiana Matejovicova, Alexandre Ramé, Morgane Rivière, et al. 2025. Gemma 3 technical report. arXiv preprint arXiv:2503.19786.
  • Wang et al. (2024) Yubo Wang, Xueguang Ma, Ge Zhang, Yuansheng Ni, Abhranil Chandra, Shiguang Guo, Weiming Ren, Aaran Arulraj, Xuan He, Ziyan Jiang, et al. 2024. MMLU-Pro: A more robust and challenging multi-task language understanding benchmark. In The Thirty-eight Conference on Neural Information Processing Systems Datasets and Benchmarks Track.
  • Yang et al. (2025) An Yang, Anfeng Li, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chang Gao, Chengen Huang, Chenxu Lv, et al. 2025. Qwen3 technical report. arXiv preprint arXiv:2505.09388.
[…]
  • (1)